# Mechanistic Modeling and Analysis of Transport Phenomena for Continuous Lyophilization

### v2

@@ Introduction
The mechanistic model presented in Section 3 is validated using the experimental data reported in Ref. [7]. Our mdel validation is carried out for each lyophilization step separately, with the specific parameters reported in Table 2.
@@
$T_{0}$ 280 K [7]
$T_{\textrm{c}},T_{\textrm{u}}$ 272 K estimated from data in Ref. [7]
$T_{\textrm{g}}$ see Fig. 5A K [7]
$T_{\textrm{n}}$ 263.18 K [7]
@@
$T_{0}$ 231 (Case 2a) K [7]
235 (Case 2b) K [7]
$T_{\textrm{b}}$ 263 (Case 2a) K [7]
313 (Case 2b) K [7]
$T_{\textrm{c}},T_{\textrm{u}}$ 275 K estimated from data in Ref. [7]
[…]
$R_{\textrm{p1}}$ 3.4$\times$107 1/s estimated from data in Ref. [7]
$R_{\textrm{p2}}$ 1 1/m estimated from data in Ref. [7]
Secondary drying
$c_{\textrm{w,0}}$ 0.088 (Case 3a) kg water/kg solid [7]
0.075 (Case 3b) kg water/kg solid [7]
$T_{0}$ 273 K [7]
$T_{\textrm{b}}$ 293 K [7]
$T_{\textrm{c}},T_{\textrm{u}}$ 285 K estimated from data in Ref. [7]
$f_{\textrm{a}}$ 0.42 1/s estimated from data in Ref. [7]
@@
$V_{\textrm{l}}$ 2$\times$10-6 m3 [7]
@@
Figure 5: Model validation for the lyophilization of suspended vials using the experimental data from Ref. [7]. Panel A shows the model prediction and experimental data for the product temperature during the freezing step. Panel B shows the model prediction and experimental data for the product temperature (assumed to be measured at the bottom surface) during the primary drying step. The maximum shelf temperatures for Cases 2a and 2b are 263 and 313 K, respectively. Panel C shows the model prediction and experimental data for the average concentration of bound water during the secondary drying step. The initial concentrations for Cases 3a and 3b are 0.088 and 0.075 kg water/kg solid, respectively.
@@
For the freezing step, the simulated temperature and freezing time agree well with the experimental data (Fig. 5A). The product temperature decreases from its initial value of 280 K to the nucleation temperature of about 263 K, where nucleation starts. Upon the first nucleation, the temperature instantaneously rises up from 263 K to the freezing point of about 272 K. The temperature slowly decreases during the solidification phase for about 30 min, following the freezing-point depression. Subsequently, the temperature decreases faster and becomes equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
@@
The model accurately predicts the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Fig. 5B). The temperature rises up rapidly at the beginning and gradually increases at a slower rate toward the end of the process. This is because the driving force for sublimation is initially low, and so most of the heat input is used to increase the temperature. As time progresses, the driving force for sublimation is larger; thus, most of the heat input is used for sublimation. The maximum error in temperature prediction is about 2–3 K, while the error in drying time prediction is about 0.2 h for Case 2a and 0.4 h for Case 2b ($\sim$6%). This error could be partly because of the uncertainty in cake resistance, which results from the stochastic nature of the freezing process.
@@
In secondary drying, the simulated concentration of bound water closely aligns with the experimental data for both concentration levels (0.088 and 0.075 kg water/kg solid). The concentration decreases exponentially from its initial value to the final concentration of 0.01 kg water/kg solid, the threshold set for terminating the simulations. This concentration profile is expected given the linear driving force model is used to describe the desorption process. It is important to note that the data used in Fig. 5C are based on the shelf temperature of 293 K. To properly estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures so that the effects of temperature on the desorption process can be quantified.
Prediction of critical process parameters for a complete continuous lyophilization cycle
@@
With the default model parameters in Table LABEL:Tab:DefaultParameters, a complete continuous lyophilization cycle can be simulated, including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Fig. 6). The time evolution of the product temperature (Fig. 6A), mass of ice (Fig. 6B), and concentration of bound water (Fig. 6C) can be obtained, given the operating pressure (Fig. 6D) and temperature (Fig. 6E). This information could be useful for various purposes. For example, it can guide process and equipment design concerning the chamber sizing and velocity profile of each vial, ensuring that the target residence time is met for each chamber. Simulation results can also be used to help identify pressure and temperature profiles to achieve some specific objectives, e.g., minimization of drying time. It is important to note that the entire simulation shown in Fig. 6 can be computed in less than 1 s on a normal laptop, which makes the model very practical to be implemented for any purposes.
@@
Visualizing spatiotemporal data
@@
Application to freezing: Optimizing freezing protocols
This section employs the model to provide insights into the freezing step, which usually receives little attention compared to the drying steps. As spontaneous nucleation was already investigated in detail by Refs. [12, 13], controlled nucleation, which is preferable for continuous lyophilization, is considered instead.
The evolution of the product temperature and amount of water during VISF at different pressures predicted by our model is shown in Fig. 9. By varying the pressure profile, it is possible to control the rate of change in the product temperature, hence nucleation. The model can also be used to estimate the amount of water that evaporates during VISF as well as the rate of evaporation. This information could be useful for the development of a freezing protocol for controlled nucleation.
Figure 9: Simulation results showing the (A) evolution of the product during temperature and (B) amount of water during VISF at three different pressures, namely $10^{4}$, 5000, and 1000 Pa. At $t=0$, the pressure decreases linearly from $10^{5}$ Pa to the target pressure in 30 s before being held constant. The nucleation temperature is set to 263 K. External heat transfer with the wall and cryogenic gas is omitted from the model so that only the contribution from VISF is observed. Other model parameters are kept at the default values.
Application to drying: Modifying the model for conventional lyophilization
Although the development of the mechanistic model in this article is for continuous lyophilization, the model can be easily modified for batch/conventional lyophilization. The major modifications are to turn off the thermal radiation part in the model and recalculate/estimate the heat transfer coefficient and cake resistance. In this case study, we consider batch primary drying and secondary drying using the data from Refs. [19] and [4], with the specific parameters reported in Table 3.
Table 3: Parameters used for the conventional drying studies, including for Cases 4a, 4b, and 5.
Symbol Value Unit Source
Primary drying
@@
$\rho_{\textrm{e}}$ 252 kg/m3 [19]
$\rho_{\textrm{f}}$ 917 kg/m3 [19]
$p_{\textrm{w,c}}$ 15 Pa [19]
$T_{0}$ 228.15 K [19]
$T_{\textrm{b}}$ see1 K [19]
$\mathcal{F}_{\textrm{s1}},\mathcal{F}_{\textrm{s3}}$ 0 omit the radiation part
@@
$R_{\textrm{p0}}$ 3.8$\times$104 m/s estimated from data in Ref. [19]
$R_{\textrm{p1}}$ 3.1$\times$107 1/s estimated from data in Ref. [19]
$R_{\textrm{p2}}$ 10 1/m estimated from data in Ref. [19]
$H$ 7.15$\times$10-3 m [19]
Secondary drying
$c_{\textrm{w,0}}$ 0.0603 kg water/kg solid [4]
$T_{0}$ 264 K [4]
$T_{\textrm{b}}$ see2 K estimated from data in Ref. [4]
$f_{\textrm{a}}$ 1.2$\times$10-3 1/s estimated from data in Ref. [4]
[…]
$\mathcal{F}_{\textrm{s1}},\mathcal{F}_{\textrm{s3}}$ 0 omit the radiation part
[…]
$V_{\textrm{l}}$ 1.5$\times$10-6 m3 [4]
$d$ 0.01425 m 2R vial, [4]
  • 1
    $T_{\textrm{b}}$ increases linearly at the rate of 0.25 K/min from 228.15 K to 258.15 K (Case 4a) and 268.15 K (Case 4b).
  • 2
    $T_{\textrm{b}}$ increases linearly at the rate of 0.5 K/min from 264 K to 312 K for Case 5.
Our model provides reliable estimates of the product temperature and sublimation interface position during primary drying at two different shelf temperatures (Fig. 10). The model can also accurately predict the product temperature and concentration of bound water during the secondary drying step (Fig. 11). With proper modifications, our continuous lyophilization model is capable of simulating conventional lyophilization.
Figure 10: Comparison of our model prediction and the experimental/simulation data from Ref. [19] for the primary drying step in conventional lyophilization. Panel A shows the model prediction and experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and reference simulation data for the sublimation front position. The maximum shelf temperatures for Cases 4a and 4b are 258.15 and 268.15 K, while other conditions are identical.
Figure 11: Comparison of our model prediction and the experimental data from Ref. [4] for the secondary drying step in conventional lyophilization. Panel A shows the model prediction and experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and experimental data for the average concentration of bound water.
Condenser failure analysis
@@
In lyophilization, one of the crucial design considerations is associated with a condenser. During normal operation, the maximum capacity of a condenser must be higher than the rate of vapor production via sublimation, and choked flow should be avoided [1]. This operational constraint is to ensure that there will not be vapor accumulation in the equipment, which could subsequently lead to in a pressure increase. This scenario may not lead to safety-related issues because the operating pressure in lyophilization is low and the amount of ice/water is generally not high enough to create overpressure. Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see (35)), which then prolongs the primary drying step. As a result, the final product quality could be affected. This section explores how to incorporate those dynamic behaviors into our model.
Most primary drying models, including our model, assume that the system pressure is well controlled, which is a reasonable assumption. To incorporate the condenser dynamics into the model, we consider cases where the total condenser capacity is not enough to condense the water vapor produced via sublimation during primary drying. In such cases, the mass balance for the water vapor in a chamber is [22]
@@
where $n_{\textrm{vial}}$ is the number of vials in the chamber. By coupling (54) and (55) with the primary drying model developed in Section 3.3, the effects of condenser failure or choked flow can be quantified.
With the specific parameters given in Table 4, the total pressure increases from the normal operating value of 3 Pa to 20 Pa (Fig. 12A). This pressure increase results from the total vapor flow exceeding the maximum condenser capacity, thus vapor accumulation in the chamber. When the system pressure increases, the driving force for sublimation (also the sublimation flux) decreases. After about 1 h, the system pressure becomes constant, indicating the equalization of the sublimation flux and condenser capacity. In this abnormal operation, the drying time and product temperature are higher than for the normal operation (Figs. 12BC).
@@
$V_{\textrm{c}}$ 0.118 m3 [22]
$j_{\textrm{w,max}}$ 1.8$\times$10-5 m3/s [22]

### v3

@@ 5.1 Model validation
The mechanistic model presented in Section 3 is validated using the experimental data reported in [Capozzi2019ContLyo_SuspendedVials]. Our model validation is carried out for each lyophilization step separately, with the specific parameters reported in Table LABEL:ch2-tab:ParametersValidation.
@@
$T_{0}$ 280 K [Capozzi2019ContLyo_SuspendedVials]
$T_{\textrm{c}},T_{\textrm{u}}$ 272 K estimated from data in [Capozzi2019ContLyo_SuspendedVials]
$T_{\textrm{g}}$ see Figure 5A K [Capozzi2019ContLyo_SuspendedVials]
$T_{\textrm{n}}$ 263.18 K [Capozzi2019ContLyo_SuspendedVials]
@@
$T_{0}$ 231 (Case 2a) K [Capozzi2019ContLyo_SuspendedVials]
235 (Case 2b) K [Capozzi2019ContLyo_SuspendedVials]
$T_{\textrm{b}}$ 263 (Case 2a) K [Capozzi2019ContLyo_SuspendedVials]
313 (Case 2b) K [Capozzi2019ContLyo_SuspendedVials]
$T_{\textrm{c}},T_{\textrm{u}}$ 275 K estimated from data in [Capozzi2019ContLyo_SuspendedVials]
@@
$R_{\textrm{p2}}$ 1 1/m estimated from data in [Capozzi2019ContLyo_SuspendedVials]
Secondary drying
$c_{\textrm{w,0}}$ 0.088 (Case 3a) kg water/kg solid [Capozzi2019ContLyo_SuspendedVials]
0.075 (Case 3b) kg water/kg solid [Capozzi2019ContLyo_SuspendedVials]
$T_{0}$ 273 K [Capozzi2019ContLyo_SuspendedVials]
$T_{\textrm{b}}$ 293 K [Capozzi2019ContLyo_SuspendedVials]
$T_{\textrm{c}},T_{\textrm{u}}$ 285 K estimated from data in [Capozzi2019ContLyo_SuspendedVials]
$f_{\textrm{a}}$ 0.42 1/s estimated from data in [Capozzi2019ContLyo_SuspendedVials]
[…]
@@
Figure 5: Model validation for the lyophilization of suspended vials using the experimental data from [Capozzi2019ContLyo_SuspendedVials]. Panel A shows the model prediction and experimental data for the product temperature during the freezing step. Panel B shows the model prediction and experimental data for the product temperature (assumed to be measured at the bottom surface) during the primary drying step. The maximum shelf temperatures for Cases 2a and 2b are 263 and 313 K, respectively. Panel C shows the model prediction and experimental data for the average concentration of bound water during the secondary drying step. The initial concentrations for Cases 3a and 3b are 0.088 and 0.075 kg water/kg solid, respectively.
@@
For the freezing step, the simulated temperature and freezing time agree well with the experimental data (Figure 5A). The product temperature decreases from its initial value of 280 K to the nucleation temperature of about 263 K, where nucleation starts. Upon the first nucleation, the temperature instantaneously rises up from 263 K to the freezing point of about 272 K. The temperature slowly decreases during the solidification phase for about 30 min, following the freezing-point depression. Subsequently, the temperature decreases faster and reaches equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
@@
The model accurately predicts the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Figure 5B). The temperature rises up rapidly at the beginning and gradually increases at a slower rate toward the end of the process. This is because the driving force for sublimation is initially low, and so most of the heat input is used to increase the temperature. As time progresses, the driving force for sublimation is larger; thus, most of the heat input is used for sublimation. The maximum error in temperature prediction is about 2–3 K, while the error in drying time prediction is about 0.2 h for Case 2a and 0.4 h for Case 2b ($\sim$6%). This error could be partly because of the uncertainty in cake resistance, which results from the stochastic nature of the freezing process. The datasets in Figure 5B, which represent the lower and upper limits of the operating boundaries, demonstrate the validity and generalizability of the model across the full range of operating conditions.
@@
Similar to primary drying, in secondary drying, we consider two different concentration levels for the bound water (residual moisture): 0.088 and 0.075 kg water/kg solid. The simulated concentration of bound water closely aligns with the experimental data for both concentration levels (Figure 5C). The concentration decreases exponentially from its initial value to the final concentration of 0.01 kg water/kg solid, the threshold set for terminating the simulations. This concentration profile is expected given the linear driving force model is used to describe the desorption process. It is important to note that the data used in Figure 5C are based on the shelf temperature of 293 K. To rigorously estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures so that the effects of temperature on the desorption process can be quantified.
Overall, our model is able to accurately predict the time evolution of the critical process parameters for all three steps of lyophilization. The maximum deviation between the predicted and measured temperature is about 3 K, which is about 1% compared to the absolute operating temperature used in lyophilization. This error is smaller than the measurement noise of some non-invasive temperature sensors, e.g., thermal imaging cameras [Srisuma2023Analytical] For the concentration of bound water, the maximum error is less than 0.01 kg water/kg solid, which is typically the threshold for identifying the endpoint of secondary drying [Fissore2018Review] Therefore, the accuracy of our model is sufficient for general process design, optimization, and control.
We note that, in the suspended-vial lyophilization considered in this article, every single vial moves through the process following the same trajectory as shown in Figure 1B, ensuring identical heat transfer conditions as a function of time for all vials. For batch lyophilization designs where the vials are distributed in different locations (Figure 1A), the evolution of the critical process parameters in each vial also varies. For instance, the outermost vials in a batch lyophilizer dry the fastest due to thermal radiation from the environment, while the thermal radiation effect is much weaker on the inner vials [Srisuma2024Rad].
5.2 Simulation of a complete lyophilization cycle
@@
With the default model parameters in Table LABEL:ch2-tab:DefaultParameters, a complete continuous lyophilization cycle can be simulated, including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Figure 6). The time evolution of the product temperature (Figure 6A), mass of ice (Figure 6B), and concentration of bound water (Figure 6C) can be obtained, given the operating pressure (Figure 6D) and temperature (Figure 6E). This information could be useful for various purposes. For example, it can guide process and equipment design concerning the chamber sizing and velocity profile of each vial, ensuring that the target residence time is met for each chamber. Simulation results can also be used to help identify pressure and temperature profiles to achieve some specific objectives, e.g., minimization of drying time. It is important to note that the entire simulation shown in Figure 6 can be computed in less than 1 s on a normal laptop, which makes the model very practical to be implemented for any purposes.
@@
5.3 Visualizing spatiotemporal data
@@
5.4 Understanding and optimizing VISF
A key benefit of mechanistic models is that they can sometimes provide important insights about the process when the experimental data are limited. This section employs the model and simulation to study the freezing step, which usually receives little attention compared to the drying steps. Spontaneous (aka uncontrolled) ice nucleation was already investigated in detail by [Deck2022FreezingLumped, Deck2024Freezing2D]. However, uncontrolled nucleation is not ideal for continuous lyophilization as described in [Capozzi2019ContLyo_SuspendedVials]. Therefore, we focus on controlled nucleation, with the VISF technique as used in [Capozzi2019ContLyo_SuspendedVials]. The input and parameter data are given in Tables 3 and 4.
Table 3: Parameters used for the modeling of VISF in Figures 9A–D.
Symbol Value Unit
$p_{\textrm{t}}$ see1 Pa
$T_{0}$ 280 K
$T_{\textrm{g}},T_{\textrm{c}},T_{\textrm{u}}$ see2 K
@@ 5.4 Understanding and optimizing VISF
  • 1
    $p_{\textrm{t}}$ is initially constant at $10^{5}$ Pa and then decreases linearly to the VISF pressure (different values are used in this study) in 1 min after VISF starts at $t$ = 0.25 h.
  • 2
    $T_{\textrm{g}},T_{\textrm{c}}$, and $T_{\textrm{u}}$ are initially constant at 268 K and then decrease linearly to 260 K in 1 min after VISF starts at $t$ = 0.25 h.
Table 4: Parameters used for the validation of our VISF model in Figure 9E.
Symbol Value Unit Source
$p_{\textrm{t}}$ see Figure 3 in [Capozzi2019ContLyo_SuspendedVials] Pa [Capozzi2019ContLyo_SuspendedVials]
$T_{0}$ 268.27 K [Capozzi2019ContLyo_SuspendedVials]
$T_{\textrm{g}}$ see Figure 3 in [Capozzi2019ContLyo_SuspendedVials] K [Capozzi2019ContLyo_SuspendedVials]
$T_{\textrm{c}},T_{\textrm{u}}$ 282 K estimated from data in [Capozzi2019ContLyo_SuspendedVials]
@@
Figure 9: Controlled nucleation via vacuum induced surface freezing (VISF). (A) Product temperature profiles at different VISF pressures when the nucleation temperature is fixed at 260 K. (B) Nucleation time and mass loss due to evaporation at different VISF pressures when the nucleation temperature is fixed at 260 K. (C) Product temperature profiles for uncontrolled nucleation and VISF at 5000 Pa and 100 Pa considering stochastic ice nucleation. (D) Nucleation time at different VISF pressures considering stochastic ice nucleation. (E) Comparison between the simulated temperature profile during VISF and experimental data from [Capozzi2019ContLyo_SuspendedVials]. Note that the nucleation time here is defined as the time when the first nucleation occurs.
Several experimental studies on VISF have been conduced in the literature. Here, we use our model and simulation results to provide insights into the process, and use that understanding to explain the previous experimental data. To clearly understand VISF, we first consider cases where that the nucleation temperature is fixed at 260 K; i.e., the effect of stochastic ice nucleation is excluded from the model (Figures 9AB). In VISF, the key idea is to reduce the system pressure, typically from the atmospheric pressure ($10^{5}$ Pa) to the VISF pressure, to evaporate a small amount of liquid, resulting in a fast decrease in the product temperature and hence nucleation. The product temperature decreases faster with a lower VISF pressure (Figure 9A). For example, when VISF starts at 0.25 h, the temperature drops from the initial value of 268 K to the nucleation temperature of 260 K almost instantaneously for the VISF pressure of 100 Pa, while it takes about 10 min when the VISF pressure is $10^{4}$ Pa. Consequently, the first nucleation occurs earlier at a lower VISF pressure (dash line in Figure 9B). During VISF, a small amount of liquid evaporates, which increases with a decrease in the VISF pressure (solid line in Figure 9B). This is expected because of a higher mass transfer driving force for evaporation at lower VISF pressure. In general, if VISF is carried out properly, the amount of liquid evaporating is small. In this case, the original liquid mass is about $2.9\times 10^{-3}$ kg, and so the mass loss of about $3\times 10^{-5}$ is almost negligible.
Next, we include the effect of stochastic ice nucleation into our VISF model for a more realistic analysis (Figures 9CD), with a Monte Carlo simulation of $10^{4}$ runs for proper statistics. Our simulation result show that VISF can significantly reduce the variation in the product temperature, compared to the uncontrolled nucleation case (Figure 9C). A decrease in the VISF pressure also reduces the degree of variation. Besides the product temperature, The variation in the nucleation time111defined as the time when the first nucleation occurs also decreases with the VISF pressure, with the first nucleation occurring earlier (Figure 9D). These results indicate that the nucleation process is well controlled with VISF (i.e., less variation in the nucleation time and temperature), agreeing with general experimental observation in the literature.
Finally, for model validation and implementation in real-world applications, the model can be fitted to experimental data obtained from the system of interest for more accurate results; the two important parameters are the mass transfer coefficient and heat transfer coefficient associated with evaporation. For example, by using the data from [Capozzi2019ContLyo_SuspendedVials], the temperature profile simulated by our model agrees well with the experimental data (Figure 9E).
The results presented in this section assume minimal vial-to-vial thermal interactions because the vials are arranged in a single row, in which all the vials experience the same heat transfer conditions throughout the process, as described in Section 3.5. This condition is not true in conventional lyophilization where a large number of vials are arranged in a hexagonal or rectangular array. In such cases, vial-to-vial interactions could significantly affect the distribution of nucleation times and freezing rates. For example, when one vial nucleates, an increase in the temperature of that vial could slightly heat the adjacent vials, which subsequently delays the first nucleation of those vials. The effects of vial-to-vial interactions on freezing have been discussed in [Deck2022FreezingLumped, Deck2024Freezing2D, Pisano2025Freezing]. In the context of modeling, these interactions can be incorporated straightforwardly by coupling our freezing model with the radiation network approach described in Section 3.5 and [Srisuma2024Rad]. In any case, such effects become much less important in controlled nucleation.
In summary, this section demonstrates how the model and simulation results elucidate the role of VISF and its operating conditions in promoting controlled nucleation and influencing the nucleation process, improving the uniformity of the product. Furthermore, we show that the model can explain real data relatively well. Hence, our model can be used to understand and guide the design and optimization of the VISF method for controlling the nucleation process in lyophilization.
5.5 Analysis of condenser failure
@@
In lyophilization, one of the crucial design considerations is associated with a condenser. During normal operation, the maximum capacity of a condenser must be higher than the rate of vapor production via sublimation, and choked flow should be avoided [Fissore2018Review]. This operational constraint is to ensure that there will not be vapor accumulation in the equipment, which could subsequently lead to in a pressure increase. This scenario may not lead to safety-related issues because the operating pressure in lyophilization is low and the amount of ice/water is generally not high enough to create overpressure. Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see Equation (35)), which then prolongs the primary drying step. As a result, the final product quality could be affected. This section explores how to incorporate those dynamic behaviors into our model.
Most primary drying models, including our model, assume that the system pressure is well controlled, which is a reasonable assumption. To incorporate the condenser dynamics into the model, we consider cases where the total condenser capacity is not enough to condense the water vapor produced via sublimation during primary drying. In such cases, the mass balance for the water vapor in a chamber is [Bano2020LumpedDrying]
@@
where $n_{\textrm{vial}}$ is the number of vials in the chamber. By coupling Equations (71) and (72) with the primary drying model developed in Section 3.3, the effects of condenser failure or choked flow can be quantified.
With the specific parameters given in Table 5, the total pressure increases from the normal operating value of 3 Pa to 20 Pa (Figure 10A). This pressure increase results from the total vapor flow exceeding the maximum condenser capacity, thus vapor accumulation in the chamber. When the system pressure increases, the driving force for sublimation (also the sublimation flux) decreases. After about 1 h, the system pressure becomes constant, indicating the equalization of the sublimation flux and condenser capacity. In this abnormal operation, the drying time and product temperature are higher than for the normal operation (Figures 10BC).
@@
$V_{\textrm{c}}$ 0.118 m3 [Bano2020LumpedDrying]
[…]